GIF++ installation and schedule meeting
Remote only
CERN
-
-
09:00
→
09:05
New Announcements 5m
AUL Test and electrical safety inspection:
Yesterday, as scheduled, we performed the AUL tests and the electrical safety visits.
Concerning the electrical safety visit, even if everything went well, some of the setups got some remarks. As soon as we receive the report, we will come back to you for fixing these minor issues.
----------------------------------------------------------------------------
Beam Pipe installation:
Beam pipe installed as scheduled on WED July 21st. It will stay till the August 11th.
Unfortunately, even if a lot of work has been performed to avoid the installation of last 2m beam pipe inside the bunker, it has been decided that we install this last part today. This installation it is crucial in order to have high purity electron beam. In case access is needed, excluding WEDs (MDs), we need to agree with LHCb (+41227663206).
Today, access to the bunker will be possible at 1pm for about 30 min.
-----------------------------------------------------------------------------Beam time July 12th to July 21st:
Even if some delay was expected, Nikos managed to give us beam on time. We started officially on Wednesday 14th July (NA FT physics) but we could have started already on Monday 12th evening.
Overall all the setups involved finished their program. Great collaboration with RD51 as reported during the SPS-PS users meeting.
Pre-Dump commissioning and RP alarm limit configuration :
Proposal : Friday 9.07 Morning
Has to be re-scheduled.----------------------------------
New radiation monitors (REMUS) :
During the maintenance week, the two REMUS detectors inside the bunker have been replaced.
The new references are :UPSTREAM one is now PMINA15402DOWNSTREAM is now PMINA15403.In addition, a new monitor has been placed before our XTDV to measure the intensity of the muon beam.
The reference for this one is PAXNA14404. Change in DIP publication done
----------------------------------
GIF pyTIMBER : Operational
1.: We have a temporary solution for the data extraction tool (with the Pytimber version that can use NXCALS). This application currently is running on a virtual machine: http://epdtdi-pytimber-gifpp.cern.ch:8080/app If you experience any issues with it, please let us know. This application is not exposed to the outside, so only accessible through the CERN GPN.
2.: The original site ( https://epdtdi-pytimber.web.cern.ch/app/gifpp ) stays deployed for the DIP/DIM live values, also accessible from outside through SSO.
----------------------------------
GIF Logbook :
Discussion ongoing. Not yet all needed features are implemented.
Not for this beam time.----------------------------------
H4 Patrol rights :
Short retraining will be needed. In case please contact us some days in advance.
---------------------------------- -
09:05
→
09:20
Weekly planning 15m
WILL BE UPDATED DURING THE MEETING !!!
----------------------------------
Weekly Planning : https://teamup.com/ksvh7dak8mub8k7xpe
UPSTREAM : Default filter settings : 2.2
DOWNSTREAM : Default filter settings : 1----------------------------------
ATLAS - MM :Yesterday July 28th the M1 in front of CMS-CSC has been reinstalled on a different trolley in vertical position. Today an access is needed for the validation of the mapping. Will profit of the stopping of the beam and the access foreseen for the installation of the beam pipe scheduled for today at 1pm.
ATLAS - sTGC :
Currently not present in bunker.ATLAS - sMDT :
The setup will stay OFF for 3 weeks. Back to the nominal position after the beam pipe is removed. To be check the gas rack, since after the power cut of yesterday it seems didn't started correctly.CMS - CSC: Some troubles restarting the DAQ after the power cuts of yesterday. Expert are working on solving the issue and as soon as the problem is solved the HV and LV will be switched ON. To be check also the battery of the CMS-CSC UPS. Yesterday also a wheel of the CMS-CSC trolley 2 broke. Intervention needed in the following days.
EP-DT2 :
Some problem with the data taking, most probably due to the electron beam passing in GIF++. Under evaluation the possibility to move the setup outside of the bunker for data taking. In case it will happen, it will be next week on Wednesday.
EcoGas: running with standard gas mixture. Two chambers where removed and transported to 904 for further tests. Back on the GIF++ bunker before September test beam.
CMS-DT: Source scan scheduled for today. After that, some decision will be taken for the attenuation to be used.CMS-RPC: Rate scan scheduled for today.
ATLAS-RPC: Giulio to confirm the installation of the A0 module. Regarding ProTOV, yesterday present in the bunker to check for the final installation.
----------Leak of SF6 suspected. We have to check on our side that it is not at the GIF++ -
09:20
→
09:30
Muon Beam Scheduling 2021 10m
Preparation and Scheduling of Muon Beam Time 2021.
Commissioning of the GIF beam dump:
Proposed date: Beginning of July (just before NA physics)
We should be ready to give feedback, e.g. counting rates etc.Week 27 - NA Setup : Beam line zones need to be CLOSED.
Nikos will contact us when we can enter the bunker.Version 1.0 of schedule for GIF & RD51 (5 weeks):
12-20 July (first 9 days)
8-14. September (1 week)
20.10 - 2.11 (2 weeks)
10-14.11 (4 days)
Planning for the first beam slot completed.CALL for RUN COORDINATORS from the systems included in the beam time !
Requirements for the control room still needs to be defined.
-
09:30
→
09:35
New requests for tests at GIF 5m
introduction, goal of the test, plan, needs
- Scintillator Irradiation for CMS HGCAL Detector.
BLM measuring campaign announced / ≈ 100 BLM, 6h approx. Week 45 (November 2nd till Friday 6th ). We will arrange everything in a way that the setup could be installed inside the bunker in the upstream area near ATL-sTGC setup -> postponed until next year
----------------------------------
Bulgaria academy of science -> testing of new type of radiation monitor systemA new setup of ALICE-RPC will come. More news in coming weeks.
Lab on Fiber (LOF) dosimeterGas recuperation system to be installed - Responsible : Roberto
4 additional racks in Gas Service area. One containing flammable gas.
FG detection system to be installed, power needs to be provided.
(DELAYED by a few months. Will be first installed in CMS, and only later come to GIF) -
09:35
→
09:40
Long Term Planning 5m
Gas usage at GIF : we need to compile a table of gas mixture / consumption
Gas training courses : ATEX level 1 and F-Gas training courses.
List of trained people to be collected.- https://lms.cern.ch/ekp/servlet/UNIVERSALSEARCH?ACTION=SEARCH&KEYW=atex
- https://lms.cern.ch/ekp/servlet/ekp?PX=N&TEACHREVIEW=N&PTX=&CID=EKP000039939&TX=FORMAT1&LANGUAGE_TAG=en&DECORATEPAGE=N
Patrol rights : Will be reset before the Muon Beam (July).
Please contact Giuseppe for re-qualification.Training Session for CESAR to be organised before beam time.
-
09:40
→
09:45
Infrastructure & Upgrades to the Facility 5m
Improvements needed or planned for the facility.
Speakers: Giuseppe Pezzullo (CERN), Martin R. Jaekel (CERN)New multi-sockets installed in the electronic racks area. Already powered and ready to use.
Monitoring of primary lines to be done. This year we will have the monitoring of the neutral gas (Ar/CO2/Helium/N2), instead next year it will be done also for the other (iC4H10, R134a, SF6, CF4 etc.). This signals will be then available in PVSS.
On going discussion for main HFO gas supply line
Removal of the fire detection interlock with the GIF++ control room. Transparent.
Ticket not closed.EN-CV: New cooling line will be installed in the preparation area.
GIFCS-Pytimber app issue still under investigation
GIFCS: new radiation monitor system (Chrome) added to our CS
-
09:45
→
09:50
General Items 5m
-
IMPACTs for 2021 5m
Guidelines for filling in IMPACT requests
List of active IMPACT declarations : https://impact.cern.ch/impact/secure/activities/accessPoint/YDALC02-PPE154
General Guideline : One IMPACT per setup.
- two chambers on individual supports -> 2 IMPACT declarations
- multiple chambers swapped on same or similar support -> one IMPACT declaration
- basically, if you need two locations, please make two IMPACT declaration (clone)
Each Setup should have IMPACT label attached with number.
Attached documents : CAD file or picture.
Review working hours : if your declaration includes the bunker area, please state the number of hours you want to stay INSIDE the bunker only (for RP purposes). -
CoVid-19 5m
CoVid-19 Risk Assessment for the Facility : https://edms.cern.ch/document/2376286/1
General guideline :
- ONE setup to enter the bunker at any given time.
- All presence at the GIF and access to the bunker needs to be announced via Teamup beforehand (!)
- Traceability of people on site : Logbook ! - List all the people you encountered.
=============================
SAFETY ALERT!
Recall of type 1 surgical masks from the DEKOTURF brand
Defects have recently been observed in DEKOTURF brand type 1 surgical masks currently distributed and in use at CERN.As a precaution, you are asked to no longer use these masks, and to return your boxes to the CERN store. Masks from another supplier (Sanitrade) are in stock and are now being distributed. To obtain the new masks, a new order must be placed following existing procedure (no exchange on site at the store for traceability reasons).Thank you for your understanding and for your collaboration. HSE Occupational Health & Safety and Environmental Protection Unit in collaboration with SCE Services & Supply Chain.
-
-
09:55
→
10:10
AOCB 15m
TRANSPORT:
Please include Francois Grenouilleau (164738) in all the EDH request related to transport.
GAS ORDERING:
- https://cern.service-now.com/service-portal?id=service_element&name=gas-support-delivery-service
Please be careful with the gas CLASS when you do an order. (please check the document here attached)
-
09:00
→
09:05